Ethanol, C2H5OH, burns in oxygen gas by the following reaction.

C2H5OH(l) + 3O2(g) → 2CO2(g) + 3H2O(g); ΔH = -1235 kJ

How many grams of ethanol would be needed to provide 293 kJ of heat?
Answer #1

when -1235.0 KJ of heat is involved, 1 mol of C2H5OH is reacting

So,

for -2.93*10^2 KJ, mol of C2H5OH = -2.93*10^2*1/-1235.0 KJ

= 0.2372 mol

Molar mass of C2H5OH,

MM = 2*MM(C) + 6*MM(H) + 1*MM(O)

= 2*12.01 + 6*1.008 + 1*16.0

= 46.068 g/mol

use:

mass of C2H5OH,

m = number of mol * molar mass

= 0.2372 mol * 46.07 g/mol

= 10.93 g

Answer: 10.9 g